OpenStudy (cocoacat):

Hi! Got stuck after first step. Write as an exact log: 2^x = 3^(x-1) 1. 2^x = 3^x * (1/3) Where should I go from here? Thanks so much!

zepdrix (zepdrix):

Noooo, let's back that step up. Take natural log as first step.

zepdrix (zepdrix):

\[\large\rm \ln(2^x)=\ln(3^{(x-1)})\]

zepdrix (zepdrix):

Apply this log rule to each side: \(\large\rm \log(a^b)=b\cdot\log(a)\)

zepdrix (zepdrix):

\[\large\rm x\ln(2)=(x-1)\ln(3)\]Then solve for x.